# Demonstrating Remote Synchronization: An Experimental Approach with Nonlinear Oscillators

### v1

@@ I Introduction
Synchronization is common in both natural and engineered systems [1, 2, 3, 4]. Studies often use globally coupled oscillator models to explore coherence [5, 6]. However, natural systems typically involve heterogeneity and local coupling, which influence behaviors such as splay states and cluster synchronization [7, 8, 9]. Remote synchronization has promising applications [10, 11, 12, 13]. In neuroscience, it offers insights into how distant brain regions coordinate without direct connections, impacting cognition and information processing [14]. In power grids, it stabilizes dispersed generators, while in communication networks, it improves data transmission and coordination, boosting performance and reliability.
@@ II-B1 Lyapunov-Floquet transformation algorithm
This section presents the Lyapunov-Floquet (Lya-Flo) transformation algorithm (1), which converts periodic linear systems into time-invariant systems for stability analysis. The algorithm is outlined step by step, detailing the necessary mathematical processes. Consider a linear time-varying system represented by the differential equation
$\dot{x}=A(t)x.$ (1)
@@ III Numerical simulation
$\dot{y}=\left[\left[\begin{array}[]{cc}0&1\\ -2\mu x_{2}x_{1}-1&\mu\left(1-x_{1}^{2}\right)\end{array}\right]-\kappa\lambda_{max}\left[\begin{array}[]{cc}1&0\\ 0&1\end{array}\right]\right]{y}$
$\dot{y}=\left[\left[\begin{array}[]{cc}0&1\\ -2\mu x_{2}x_{1}-1&\mu\left(1-x_{1}^{2}\right)\end{array}\right]+\gamma\left[\begin{array}[]{cc}1&0\\ 0&1\end{array}\right]\right]{y}$ (7)
@@ IV Experimental results
Figure 7: Experimental validation of synchronization in a coupled VPO network with arbitrary topology. The x-axis shows time (s), while the y-axis shows the node voltages of x1 and y1 from the two clusters.

### v2

@@ I Introduction
Synchronization is common in both natural and engineered systems [1, 2, 3, 4]. Studies often use globally coupled oscillator models to explore coherence [5, 6, 7]. However, natural systems typically involve heterogeneity and local coupling, which influence behaviors such as splay states and cluster synchronization [8, 9, 10]. Remote synchronization has promising applications [11, 12, 13, 14]. In neuroscience, it offers insights into how distant brain regions coordinate without direct connections, impacting cognition and information processing [15, 16]. In power grids, it stabilizes dispersed generators, while in communication networks, it improves data transmission and coordination, boosting performance and reliability.
@@ II-B1 Lyapunov-Floquet transformation algorithm
This section presents the Lyapunov-Floquet (Lya-Flo) transformation algorithm (1),
@@
which converts periodic linear systems into time-invariant systems for stability analysis [18]. The algorithm is outlined step by step, detailing the necessary mathematical processes.
@@ III Numerical simulation
$\dot{\eta}=\left[\left[\begin{array}[]{cc}0&1\\ -2\mu x_{2}x_{1}-1&\mu\left(1-x_{1}^{2}\right)\end{array}\right]-\kappa\lambda_{max}\left[\begin{array}[]{cc}1&0\\ 0&1\end{array}\right]\right]{\eta}$
$\dot{\eta}=\left[\left[\begin{array}[]{cc}0&1\\ -2\mu x_{2}x_{1}-1&\mu\left(1-x_{1}^{2}\right)\end{array}\right]+\gamma\left[\begin{array}[]{cc}1&0\\ 0&1\end{array}\right]\right]{\eta}$ (6)
@@ IV Experimental results
Figure 7: Experimental validation of synchronization in a coupled Van der Pol oscillator network with arbitrary topology. The x-axis shows time (s), while the y-axis shows the node voltages of x1 and y1 from the two clusters.